the legs 12 cm and 5 cm what the hypotenuse
Answer by mananth(16946)

By pythagoras theorem
side1^2+side2^2= hypotenuse ^2
12^2+5^2= hypotenuse^2
144+25=hypotenuse ^2
169 = hyptenuse ^2
square root
hypotenuse = 13 cm